In a group of 125 computer users, 50 of them have a GPU installed in their systems, 30 of them have SSD storage, and 15 have both. If a computer user chosen at random has a GPU, what is the probability he/she also has an SSD?


1
Expert's answer
2021-04-19T06:40:28-0400


Let the events:

P(A) = probability that a computer has a GPU installed = 50/120 = 0.4

P(B) = probability that a computer has SSD = 30/120 = 0.24

and

P(A and B) = 15/125 = 0.12

The two events are not independent since;


P(A)*P(B) = 0.4 X 0.24 = 0.096, which is and not equal to P(A and B)

For dependent events:

P(B|A) = P(A and B)/P(A)

P(B|A) = 0.12/0.4 = 0.3

Therefore, Β if a computer user chosen at random has a GPU, the probability he/she also has an SSD = 0.3 or 30%